at constant temperature the pressure on a 4.5 l sample of gas is increased from 1.2 atm to 2.3 atm. what is the new volume of the gas sample


1
Expert's answer
2020-09-29T06:56:20-0400

According to Boyle's law "P1V1= P2V2"

So here at constant temperature and for a constant amount of gass we can use this to find out the final volume V2

P1= 1.2 atm

P2= 2.3 atm

V2 = ?

V1 = 4.5 lit

So V2 = (P1×V1) /P2

= (1.2 ×4.5) /2.3

= 5.4/2.3

= 2.35 Lit is ans. This is the new volume of the gass